A 100 g piece of ice at initial temperature 0 C is added to 1 L water at initial temperature 100 C. What is the final temperature of the system?



Please show all steps and work . Thank you very much for help!
Answer #1

heat gained by the ice Q loss = mice*Lf + mice*Sw*dt1

Qloss = 0.1*334000 + (0.1*4190*t) = 33400 + 419t

heat lost by water Q gain = mwater*sw*dt2 = d_water*volume*sw*(100-t)

Qloss = 1000*1e-3*4190*(100-t) = 419000 - 4190 t

Qloss = Qgain


419000 - 4190 t = 33400 + 419t


t = 83.66 oC